# On-the-Fly Adaptation to Quantization: Configuration-Aware LoRA for Efficient Fine-Tuning of Quantized LLMs

### v1

@@ 1 Introduction
With the rapid growth of parameter scale, large language models (LLMs) have demonstrated increasingly strong capabilities across a wide range of applications (Zhang et al., 2022; Touvron et al., 2023; Liu et al., 2024a; Workshop et al., 2022). However, their model size makes their deployment on edge devices impractical. A common solution is to first quantize the pretrained model and then fine-tune it using Low-Rank Adaptation (LoRA) (Hu et al., 2022), allowing model compression while maintaining performance (Dettmers et al., 2023; Guo et al., 2024; Xu et al., 2024). A key factor in this process is the quantization configuration, which specifies the bit-widths of layers in the model and thereby determines the overall compression level. However, existing methods are typically designed for a single fixed configuration and thus cannot generalize across diverse configuration settings. In practice, this limitation becomes critical because edge devices, ranging from smartphones to laptops, require support for diverse compression levels. Consequently, a single LoRA adapter is insufficient to deliver consistent performance across all possible configurations. Although fine-tuning a separate LoRA adapter for each configuration is possible, this can be extremely time-consuming.
In this work, we aim to address the challenge of adjusting LoRA adapters to arbitrary quantization configurations in an efficient manner. This is achieved through a configuration-aware model that learns to associate each configuration with its corresponding adapter adjustments. This goal raises two key challenges. First, directly mapping from each configuration to the full set of LoRA parameters is intractable, as the output space of the mapping is prohibitively large. Second, the effectiveness of this model hinges on the quality of the training configuration set. Yet, constructing such a configuration set is nontrivial: uniform bit-width assignments typically fail to account for the heterogeneous sensitivity of different layers, resulting in suboptimal results. To address these challenges, we propose CoA-LoRA, a configuration-aware model that integrates two key techniques: (i) a lightweight mapping from each layer’s quantization configuration to a compact low-rank adjustment, which effectively reduces the dimensionality of the mapping output and enables parallel adjustment across layers; and (ii) a configuration set search based on a Pareto-based Gaussian process (Williams & Rasmussen, 1995), which refines the training configuration set and guides more accurate LoRA adjustment. By integrating these two techniques, CoA-LoRA produces high-quality LoRA adapters that generalize well across heterogeneous devices, eliminating the need for repeated fine-tuning when encountering new configurations.
@@
  • Empirical experiments show that CoA-LoRA efficiently serves all quantization configurations with a single trained configuration-aware model, avoids the need for separate fine-tuning of each configuration, and achieves superior performance on most datasets with accuracy gains ranging from $2.36\%$ to $10.37\%$ over state-of-the-art methods.
@@ 2.2 Weight Quantization of Large Language Models
Weight quantization, which quantizes the model weights while keeping activations in full precision, is a widely used and practical approach for deploying LLMs under memory and compute constraints. Standard round-to-nearest (RTN) quantization (Yao et al., 2021) is one of the most widely used techniques, where weights $\bm{w}$ are quantized as $\bm{w}\approx s\times\text{clamp}\!\left(\left\lfloor\frac{\bm{w}}{s}\right\rceil;-2^{b-1},2^{b-1}-1\right)$ with scaling factor $s=\frac{\max(\bm{w})}{2^{b-1}-1}$ and bit-width $b$. NormalFloat (NF) (Dettmers et al., 2022) offers an alternative by preserving the floating-point representation with reduced exponent and fractional bits, allowing better handling of extreme values under aggressive low-bit quantization. Despite their advantages, both RTN and NF suffer from substantial accuracy degradation at very low bit-widths (Xiao et al., 2023; Shao et al., 2024). To alleviate this, several advanced low-bit quantization methods have been proposed. For example, GPTQ (Frantar et al., 2023) leverages second-order information of the quantization error for precise 3- or 4-bit weight quantization, while AWQ (Lin et al., 2024) and OWQ (Lee et al., 2024) allocate quantization precision based on weight importance. Nevertheless, even with these techniques, quantized LLMs typically experience performance drops compared to pretrained LLMs. Q-LoRA (Dettmers et al., 2023) addresses this issue by applying LoRA technique to fine-tune quantized weights, partially restoring performance. Building on this idea, LQ-LoRA (Guo et al., 2024) further restores performance by initializing the LoRA adapter using iterative singular value decomposition (SVD), which provides a better starting point for fine-tuning. However, in realistic scenarios involving large-scale and heterogeneous devices, performing LoRA fine-tuning for every quantization configuration becomes computationally infeasible. This limitation motivates approaches that can adjust a single LoRA adapter across multiple quantization settings, enabling efficient deployment of quantized LLMs without repeated fine-tuning.
@@ 3 Motivation
different quantization configurations. To illustrate these challenges, Fig. 1 compares Q-LoRA and Shared-LoRA on the SST-2 task (Wang et al., 2019), using RoBERTa-Large (Liu et al., 2019). Q-LoRA fine-tunes a dedicated adapter for each quantization setting (2.5–4 bits), yielding strong accuracy but requiring fine-tuning effort that grows linearly with the number of quantized models. Shared-LoRA eliminates this cost by training a shared adapter across all settings, but the gain in efficiency comes with a large drop in accuracy.
These observations naturally motivate the following question: can we design a method that efficiently adjust a LoRA adapter across different quantization settings without repeated fine-tuning?
@@ 4.1 Overview of CoA-LoRA
These two techniques are executed in a cyclic training process: (i) Train the configuration-aware model $\bm{\theta}$ on the current configuration set $\mathcal{C}$; (ii) Expand $\mathcal{C}$ through gradient-guided search, and refine $\mathcal{C}$ via diversity-preserving Pareto filtering. Once trained, $\bm{\theta}$ can efficiently generate LoRA adjustments for any configuration, thus avoiding repeated fine-tuning.
@@ 4.2 Configuration-Aware LoRA Adjustment
Quantization Configuration Representation. We adopt the NormalFloat (NF) quantization scheme (Dettmers et al., 2023) as the quantization method to determine a quantization configuration. NF is particularly suitable in our setting because it supports block-wise quantization at a finer granularity, which enables a more expressive configuration space.
@@
Under the corresponding layer-level configuration $\bm{c}_{i}$, layer $i$ contains a pair of low-rank matrices $\mathbf{L}_{1,i},\mathbf{L}_{2,i}$} that need to be adjusted. According to the layer-level configuration space of parameters listed in Table 1, the overall search space is $(4\cdot 4\cdot 3\cdot 3\cdot 3)^{N}$, which grows exponentially with $N$. To make this space tractable for learning, we embed each layer-level configuration $\bm{c}_{i}$ into a learned vector $\bm{z}_{i}$. In addition, we further embed the layer name (e.g., fully connected layers or attention projection layers) and the block index into vectors $\bm{m}$ and $\bm{b}$, respectively. The final embedding for layer $i$ under the $j$-th model-level configuration is given by $\mathbf{C}_{i}^{(j)}\triangleq[\bm{z}_{i}^{(j)},\bm{m}_{i}^{(j)},\bm{b}_{i}^{(j)}]$. Based on these layer embeddings, the $j$-th model-level configuration embedding is defined as their concatenation $\mathbf{C}^{(j)}=[\mathbf{C}_{1}^{(j)},\dots,\mathbf{C}_{N}^{(j)}]$.
@@
Motivated by this observation, we introduce a configuration-aware model $\bm{\theta}:\mathbb{R}^{|\mathbf{C}_{i}|}\rightarrow\mathbb{R}^{r\times r}$, which maps a layer-level configuration vector $\mathbf{C}_{i}$ to a lightweight adjustment matrix $\mathbf{U}_{\bm{\theta}}(\mathbf{C}_{i})\in\mathbb{R}^{r\times r}$. As shown in Fig. 3, each layer’s low-rank matrix $\mathbf{L}_{2,i}$ is reparameterized as $(\mathbf{I}+\mathbf{U}_{\bm{\theta}}(\mathbf{C}_{i}))\mathbf{L}_{2,i}$, where $\mathbf{I}$ is the identity matrix. Given a dataset $\mathcal{D}$, let $\widetilde{\mathbf{W}}_{\mathbf{C}}$ denote the quantized pre-trained model weights under configuration $\mathbf{C}=[\mathbf{C}_{1},...,\mathbf{C}_{N}]$. We define the adjusted model weights using a configuration-aware adjustment function:
$\widetilde{\mathbf{W}}_{\mathbf{C}}^{\text{LoRA}}=\text{InsertLoRA}\Big(\widetilde{\mathbf{W}}_{\mathbf{C}},\big\{\mathbf{L}_{1,i}(\mathbf{I}+\mathbf{U}_{\bm{\theta}}(\mathbf{C}_{i}))\mathbf{L}_{2,i}\big\}_{i=1}^{N}\Big),$ (2)
where $\text{InsertLoRA}(\cdot)$ inserts each layer’s LoRA adjustment into the corresponding layer.
@@ 4.3 Pareto-Based Quantization Configuration Search and Filtering
To improve the quality of the initialized configuration set $\mathcal{C}$ for training the configuration-aware model, we iteratively update it through a Pareto-based search. Each candidate configuration is evaluated according to two inherently conflicting objectives: task-specific performance $f_{1}$ and average bit-width $f_{2}$, as higher performance typically requires higher precision. We therefore formulate a bi-objective optimization problem for each configuration $\mathbf{C}$:
@@
By identifying the Pareto-optimal configurations—those for which no other configuration improves one objective without degrading the other—we obtain a set of configurations that is both high-performing and diverse, forming the selected training configuration set $\mathcal{C}$. Projecting these configurations onto the objective space $(f_{1},f_{2})$ defines the Pareto front. To evaluate the quality of a set of trade-off configurations, we use the hypervolume (HV) metric (Zitzler & Thiele, 1999). Given a reference point $\bm{r}$, the hypervolume $\mathcal{H}_{\bm{r}}$ measures the area dominated by the Pareto front. In Fig. 4 (left), the combined green and yellow areas indicate the hypervolume formed by the five trade-off configurations with respect to $\bm{r}$, where a larger HV value indicates a better Pareto front.
@@
where $m(\mathbf{C})$ is the mean function, $k(\mathbf{C},\mathbf{C}^{\prime})$ is a kernel (e.g., RBF), and $\mathbf{C},\mathbf{C}^{\prime}\in\mathcal{C}$. To select a configuration that maximally contributes to the hypervolume of the current configuration set, we use the Expected Hypervolume Improvement (EHVI):
@@
$\mathcal{C}^{(u)}_{\mathrm{Pareto}}=\{\mathbf{C}\in\mathcal{C}_{u}\mid\mathbf{f}(\mathbf{C})\nsucc\mathbf{f}(\mathbf{C}^{\prime})\ \text{for all }\mathbf{C}^{\prime}\in\mathcal{C}_{u},\mathbf{C}^{\prime}\neq\mathbf{C}\},$ (9)
where $\mathbf{f}(\mathbf{C})\nsucc\mathbf{f}(\mathbf{C}^{\prime})$ means that $\mathbf{C}$ is not dominated by $\mathbf{C}^{\prime}$, i.e., there exists at least one objective for which $\mathbf{f}(\mathbf{C})$ is not worse than $\mathbf{f}(\mathbf{C}^{\prime})$.
@@ 5.1 Experimental Setting
We evaluate CoA-LoRA in two settings: (I) language modeling fine-tuning on the C4 dataset (Dodge et al., 2021), and (II) fine-tuning on several GLUE tasks, namely QNLI, MNLI, SST-2, and QQP (Wang et al., 2019). For setting (I), we use Qwen-2.5-1.5B, Qwen-2.5-3B (Qwen et al., 2025), and LLaMA-2-7B (Touvron et al., 2023), whereas for setting (II), we employ RoBERTa-Large model (Liu et al., 2019).
Baselines. We compare CoA-LoRA with several baselines: Q-LoRA (Dettmers et al., 2023), which quantizes the pretrained model before LoRA fine-tuning; LQ-LoRA (Guo et al., 2024), which quantizes the pretrained model and initializes LoRA with SVD; GPTQ-LoRA, which applies GPTQ quantization followed by LoRA fine-tuning; and Shared-LoRA, which fine-tunes multiple quantized models using a single LoRA adapter. Among them, LQ-LoRA serves as the strongest baseline due to its near-optimal initialization and one-to-one fine-tuning. Our goal is for CoA-LoRA to achieve performance comparable to LQ-LoRA and Q-LoRA, while avoiding repeated fine-tuning by efficiently adjusting LoRA adapters to new quantization settings.
Evaluation. Following prior work (Guo et al., 2024; Dettmers et al., 2023), we evaluate model performance across GLUE’s benchmark tasks using accuracy. For models fine-tuned on C4, we report perplexity on the validation set. To demonstrate the effectiveness of CoA-LoRA in efficiently adapting low-rank matrices across arbitrary bit-widths, we present its performance over 50 configurations spanning 2.5 to 6.25 bits. For Q-LoRA and LQ-LoRA, due to the high computational cost of fine-tuning—each new configuration requires a separate fine-tuning run—we only report results for 2.5, 3, 3.5, and 4 bits. We use hypervolume (HV) (Zitzler & Thiele, 1999) to evaluate the performance curves generated by each algorithm for different bit-widths. A larger HV indicates that the algorithm achieves higher performance while covering a broader range of bit-widths. Additionally, we report the total training time of each algorithm, as well as the average accuracy gap relative to Q-LoRA computed over four bit-widths.
@@
Table 2: Comparison of HV, accuracy gap, and training time across four tasks. Training Times are reported in minutes (m). The best HV and accuracy gap per task are highlighted in bold. The accuracy gap is computed relative to Q-LoRA, and the dash (“-”) indicates the baseline itself.
@@
Q-LoRA 4 0.57 - 79m 0.54 - 132m 0.63 - 65m 0.52 - 126m
LQ-LoRA 4 0.59 +6.22% 72m 0.56 +5.95% 122m 0.64 +2.84% 61m 0.52 +2.68% 115m
Shared-LoRA 1 0.45 -15.93% 21m 0.57 +8.93% 35m 0.61 -6.14% 19m 0.53 +1.31% 32m

### v2

@@ 1 Introduction
With the rapid growth of parameter scale, large language models (LLMs) have demonstrated increasingly strong capabilities across a wide range of applications (Zhang et al., 2022; Touvron et al., 2023; Liu et al., 2024a; Workshop et al., 2022; Ye and Tang, 2025). Nevertheless, their model size makes deployment on edge devices impractical. A common solution is to first quantize the pretrained model and then fine-tune it using Low-Rank Adaptation (LoRA) (Hu et al., 2022), allowing model compression while maintaining performance (Dettmers et al., 2023; Guo et al., 2024; Xu et al., 2024). A key factor in this process is the quantization configuration, which specifies the bit-widths of layers in the model and thereby determines the overall compression level. However, existing methods are typically designed for a single fixed configuration and thus cannot generalize across diverse configuration settings. In practice, this limitation becomes critical because edge devices, ranging from smartphones to laptops, require support for diverse compression levels. Consequently, a single LoRA adapter is insufficient to deliver consistent performance across all possible configurations. Although fine-tuning a separate LoRA adapter for each configuration is possible, this can be extremely time-consuming.
In this work, we aim to address the challenge of adjusting LoRA adapters to arbitrary quantization configurations in an efficient manner. This is achieved through a configuration-aware model that learns to associate each configuration with its corresponding adapter adjustments. This goal raises two key challenges. First, directly mapping from each configuration to the full set of LoRA parameters is intractable, as the output space of the mapping is prohibitively large. Second, the effectiveness of this model hinges on the quality of the training configuration set. Yet, constructing such a configuration set is nontrivial: uniform bit-width assignments typically fail to account for the heterogeneous sensitivity of different layers, resulting in suboptimal results. To address these challenges, we propose CoA-LoRA, a configuration-aware model that integrates two key techniques: (i) a lightweight mapping from each layer’s quantization configuration to a compact low-rank adjustment, which effectively reduces the dimensionality of the mapping output and enables parallel adjustment across layers; and (ii) a configuration set search based on a Pareto-based Gaussian process (Williams and Rasmussen, 1995), which refines the training configuration set and guides more accurate LoRA adjustment. By integrating these two techniques, CoA-LoRA produces high-quality LoRA adapters that generalize well across heterogeneous devices, eliminating the need for repeated fine-tuning when encountering new configurations.
@@
  • Empirical experiments show that CoA-LoRA efficiently serves all quantization configurations with a single trained configuration-aware model, avoids the need for separate fine-tuning of each configuration, and achieves superior performance on most datasets with accuracy gains ranging from $1.74\%$ to $8.89\%$ over state-of-the-art methods.
@@ 2.2 Weight Quantization of Large Language Models
Weight quantization, which quantizes the model weights while keeping activations in full precision, is a widely used and practical approach for deploying LLMs under memory and compute constraints. Standard round-to-nearest (RTN) quantization (Yao et al., 2021) is one of the most widely used techniques, where weights $\bm{w}$ are quantized as $\bm{w}\approx s\times\text{clamp}\!\left(\left\lfloor\frac{\bm{w}}{s}\right\rceil;-2^{b-1},2^{b-1}-1\right)$ with scaling factor $s=\frac{\max(\bm{w})}{2^{b-1}-1}$ and bit-width $b$. NormalFloat (NF) (Dettmers et al., 2023) is a 4-bit floating-point format that employs non-uniform quantization via a lookup table to approximate the original floating-point values. Despite their advantages, RTN may suffer noticeable performance drop at extremely low bit-widths (Shao et al., 2024); NF initially employed uniform granularity, assigning the same bitwidth to every layer, which restricted fine-grained layer-wise adjustments and limited potential performance gains (Zhou et al., 2025). To alleviate this, several advanced low-bit quantization methods have been proposed. For example, GPTQ (Frantar et al., 2023) leverages second-order information of the quantization error for precise 3- or 4-bit weight quantization, while AWQ (Lin et al., 2024) and OWQ (Lee et al., 2024) allocate quantization precision based on weight importance. Nevertheless, even with these techniques, quantized LLMs typically experience performance drops compared to pretrained LLMs. QLoRA (Dettmers et al., 2023) addresses this issue by applying LoRA technique to fine-tune quantized weights, partially restoring performance. Building on this idea, LQ-LoRA (Guo et al., 2024) further restores performance by initializing the LoRA adapter using iterative singular value decomposition (SVD), which provides a better starting point for fine-tuning. However, in realistic scenarios involving large-scale and heterogeneous devices, performing LoRA fine-tuning for every quantization configuration becomes computationally infeasible. This limitation motivates approaches that can adjust a single LoRA adapter across multiple quantization settings, enabling efficient deployment of quantized LLMs without repeated fine-tuning.
2.3 LoRA Adapter Generation
Early works such as P-diff families (Wang et al., 2024) and G.pt (Peebles et al., 2022) pioneered the use of diffusion models to generate network parameters. More recently, CONDP-DIFF (Jin et al., 2024) applies diffusion models to generate LoRA parameters in multi-task learning scenarios. While effective for coarse-grained tasks, it struggles with fine-grained tasks. To address this limitation, ICM-LoRA (Shao et al., 2025) leverages in-context learning, and LoRA-Gen (Xiao et al., 2025) improves both efficiency and performance using Mixture-of-Experts (MoE) (Lepikhin et al., 2021). However, these methods generally rely on expert LoRA and well-trained LoRA parameters, which are costly to obtain for quantized LLMs. In contrast, our work aims to adjust LoRA adapters in a lightweight and effective manner, restoring model performance under different quantization settings.
@@ 3 Motivation
different quantization configurations. To illustrate these challenges, Fig. 1 compares QLoRA and Shared-LoRA on the SST-2 task (Wang et al., 2019), using RoBERTa-Large (Liu et al., 2019). QLoRA fine-tunes a dedicated adapter for each quantization setting (2.5–4 bits), yielding strong accuracy but requiring fine-tuning effort that grows linearly with the number of quantized models. Shared-LoRA eliminates this cost by training a shared adapter across all settings, but the gain in efficiency comes with a large drop in accuracy.
These observations naturally motivate the following question: can we design a method that efficiently adjust the LoRA adapter across different quantization settings without repeated fine-tuning?
@@ 4.2 Configuration-Aware LoRA Adjustment
Quantization Configuration Representation. We adopt the NormalFloat (NF) quantization scheme (Dettmers et al., 2023) as the quantization method to determine a quantization configuration.
@@
We choose NormalFloat (NF) primarily because its non-uniform quantization allows better preservation of small-magnitude weights compared to uniform integer-based quantization at the same bit-width, which is critical for maintaining performance in low-bit quantization.
@@
Under the corresponding layer-level configuration $\bm{c}_{i}$, layer $i$ contains a pair of low-rank matrices $\mathbf{L}_{1,i},\mathbf{L}_{2,i}$} that need to be adjusted. According to the layer-level configuration space of parameters listed in Table 1, the overall search space is $(4\cdot 4\cdot 3\cdot 3\cdot 3)^{N}$, which grows exponentially with $N$. To make this space tractable for learning, we embed each layer-level configuration $\bm{c}_{i}$ into a learned vector $\bm{z}_{i}$. In addition, we further embed the layer name (e.g., fully connected layers or attention projection layers) and the block index into vectors $\bm{m}$ and $\bm{b}$, respectively. The final embedding for layer $i$ under the $j$-th model-level configuration is given by $\mathbf{Q}_{i}^{(j)}\triangleq[\bm{z}_{i}^{(j)},\bm{m}_{i}^{(j)},\bm{b}_{i}^{(j)}]$. Based on these layer embeddings, the $j$-th model-level configuration continuous embedding is defined as their concatenation $\mathbf{Q}^{(j)}=[\mathbf{Q}_{1}^{(j)},\dots,\mathbf{Q}_{N}^{(j)}]$. In the following, we denote the corresponding discrete quantization configuration as $\mathbf{C}^{(j)}$.
@@
Motivated by this observation, we introduce a configuration-aware model $\bm{\theta}:\mathbb{R}^{|\mathbf{Q}_{i}|}\rightarrow\mathbb{R}^{r\times r}$, which maps a layer-level configuration vector $\mathbf{Q}_{i}$ to a lightweight adjustment matrix $\mathbf{U}_{\bm{\theta}}(\mathbf{Q}_{i})\in\mathbb{R}^{r\times r}$. As shown in Fig. 3, each layer’s low-rank matrix $\mathbf{L}_{2,i}$ is reparameterized as $(\mathbf{I}+\mathbf{U}_{\bm{\theta}}(\mathbf{Q}_{i}))\mathbf{L}_{2,i}$, where $\mathbf{I}$ is the identity matrix. Given a dataset $\mathcal{D}$, let $\widetilde{\mathbf{W}}_{\mathbf{C}}$ denote the quantized pre-trained model weights under configuration $\mathbf{C}$. We define the adjusted model weights using a configuration-aware adjustment function:
$\widetilde{\mathbf{W}}_{\mathbf{C}}^{\text{LoRA}}=\text{InsertLoRA}\Big(\widetilde{\mathbf{W}}_{\mathbf{C}},\big\{\mathbf{L}_{1,i}^{(\mathbf{C})}(\mathbf{I}+\mathbf{U}_{\bm{\theta}}(\mathbf{Q}_{i}))\mathbf{L}_{2,i}^{(\mathbf{C})}\big\}_{i=1}^{N}\Big),$ (2)
where $\text{InsertLoRA}(\cdot)$ inserts each layer’s LoRA adjustment into the corresponding layer. During the generation of the LoRA adjustment, our algorithm computes the residual between pretrained weights and the quantized weights under $\mathbf{C}$ and applies an SVD to derive $\mathbf{L}_{1,i}^{(\mathbf{C})}$ and $\mathbf{L}_{2,i}^{(\mathbf{C})}$.
@@ 4.3 Pareto-Based Quantization Configuration Search and Filtering
To improve the quality of the initialized configuration set $\mathcal{Q}$ for training the configuration-aware model, we iteratively update it through a Pareto-based search. Each candidate configuration is evaluated according to two inherently conflicting objectives: task-specific performance $f_{1}$ and average bit-width $f_{2}$, as higher performance typically requires higher precision. We therefore formulate a bi-objective optimization problem for each configuration $\mathbf{C}$:
@@
The computation of $f_{2}$ is provided in the Appendix (Eq. 12). By identifying the Pareto-optimal configurations—those for which no other configuration improves one objective without degrading the other—we obtain a set of configurations that is both high-performing and diverse, forming the selected training configuration set $\mathcal{C}$. Projecting these configurations onto the objective space $(f_{1},f_{2})$ defines the Pareto front. To evaluate the quality of a set of trade-off configurations, we use the hypervolume (HV) metric (Zitzler and Thiele, 1999). Given a reference point $\bm{r}$, the hypervolume $\mathcal{H}_{\bm{r}}$ measures the area dominated by the Pareto front. In Fig. 4 (left), the combined green and yellow areas indicate the hypervolume formed by the five trade-off configurations with respect to $\bm{r}$, where a larger HV value indicates a better Pareto front.
@@
where $m(\mathbf{C})$ is the mean function, $k(\mathbf{C},\mathbf{C}^{\prime})$ is a kernel (e.g., RBF), and $\mathbf{C},\mathbf{C}^{\prime}\in\mathcal{C}$.
To select a configuration that maximally contributes to the hypervolume of the current configuration set, we use the Expected Hypervolume Improvement (EHVI):
@@
$\mathcal{C}^{(u)}_{\mathrm{Pareto}}=\{\mathbf{C}\in\mathcal{C}_{u}\mid\mathbf{f}(\mathbf{C}^{\prime})\nsucc\mathbf{f}(\mathbf{C})\ \text{for all }\mathbf{C}^{\prime}\in\mathcal{C}_{u},\mathbf{C}^{\prime}\neq\mathbf{C}\},$ (9)
where $\mathbf{f}(\mathbf{C}^{\prime})\nsucc\mathbf{f}(\mathbf{C})$ means that $\mathbf{C}$ is not dominated by $\mathbf{C}^{\prime}$, i.e., there exists at least one objective for which $\mathbf{f}(\mathbf{C})$ is not worse than $\mathbf{f}(\mathbf{C}^{\prime})$.
@@ 5.1 Experimental Setting
We evaluate CoA-LoRA in two settings: (I) fine-tuning on the C4 dataset (Dodge et al., 2021), and (II) fine-tuning on several GLUE tasks, namely QNLI, MNLI, SST-2, and QQP (Wang et al., 2019). For setting (I), we use Qwen-2.5-1.5B, Qwen-2.5-3B (Qwen et al., 2025), and LLaMA-2-7B (Touvron et al., 2023), whereas for setting (II), we employ RoBERTa-Large model (Liu et al., 2019). Our implementation is available at https://github.com/rG223/CoA-LoRA.
Baselines. We compare CoA-LoRA with several baselines: QLoRA (Dettmers et al., 2023), which quantizes the pretrained model before LoRA fine-tuning; LQ-LoRA (Guo et al., 2024), which quantizes the pretrained model and initializes LoRA with SVD; GPTQLoRA, which applies GPTQ quantization followed by LoRA fine-tuning; and Shared-LoRA, which fine-tunes multiple quantized models using a single LoRA adapter. Among them, LQ-LoRA serves as the strongest baseline due to its near-optimal initialization and one-to-one fine-tuning. Our goal is for CoA-LoRA to achieve performance comparable to LQ-LoRA and QLoRA, while avoiding repeated fine-tuning by efficiently adjusting LoRA adapters to new quantization settings.
Evaluation. Following prior work (Guo et al., 2024; Dettmers et al., 2023), we evaluate model performance across GLUE’s benchmark tasks using accuracy. For models fine-tuned on C4, we report perplexity on the validation set. For a fair comparison, all baseline methods are quantized in a layer-wise manner; given an average bit-width across layers, the corresponding layer-wise quantization configurations are automatically determined (see Eq. (11) in the appendix). To demonstrate the effectiveness of CoA-LoRA, we present its performance over 50 configurations spanning 2.5 to 6.25 bits. For QLoRA and LQ-LoRA, due to the high computational cost of fine-tuning—each new configuration requires a separate fine-tuning run—we only report results for 2.5, 3, 3.5, 4, and 6 bits. We use hypervolume (HV) (Zitzler and Thiele, 1999) to evaluate the performance curves generated by each algorithm for different bit-widths. A larger HV indicates that the algorithm achieves higher performance while covering a broader range of bit-widths. Additionally, we report the total training time of each algorithm, as well as the average accuracy gap relative to QLoRA computed over four bit-widths.
@@
Additional experiments are provided in Appendix C, including the impact of the number of segments $U$ and zero-shot accuracy comparisons across downstream tasks, as well as other visualizations.
Table 2: Comparison of HV, accuracy gap, and training time across four tasks. Training Times are reported in minutes (m). The best HV and accuracy gap per task are highlighted in bold. The average accuracy gap is computed relative to QLoRA, and the dash (“-”) indicates the baseline itself.
@@
QLoRA 6 0.58 - 119m 0.54 - 208m 0.63 - 97m 0.54 - 189m
LQ-LoRA 6 0.59 +2.81% 108m 0.57 +8.13% 183m 0.64 +1.54% 91m 0.54 +0.47% 172m
Shared-LoRA 1 0.60 +2.90% 21m 0.57 +8.11% 35m 0.61 -5.06% 19m 0.53 -1.18% 32m